a certain sum of money doubles itself at simple interest in 8 years. in how many years will it be three times at the same rate? Answer by mananth(16949) 
 principal =x
Interest =x
x= x*8*r/100
r=100/8
--------------=
2x= x*n*100/8 *1/100
2=100n/800
1600=100n
n= 16 years